# NSD-DIL: Null-Shot Deblurring Using Deep Identity Learning

### v1

@@ 4.7 Conclusion
We have established a new problem formulation based on the DIL objective for the deblurring task, which can be extended to the ISR task as well. The proposed NSD-DIL is the first DL-based image data-independent deblurring model that involves a restoration process rather than learning mapping from distorted to sharp image distribution. The proposed method uses the degradation kernel space effectively to restore the sharp images from blurred images based on the proposed dedicated regularization term provided in the DIL objective function. Our NSD-DIL method paves the path to have a deeper look into the learning and understanding of the inverse degradation kernel since the explicit representation i.e. DRK, in a matrix form is made possible. Additionally, we have demonstrated the deblurring task of replacing the deep model with a simple convolution operation with the computed DRK, which doesn’t require high-performance computing resources. The proposed method’s performance is superior to the SotA works, despite being a very lightweight deblurring model. The experimental results and computational performance comparisons with the SotA indicate that the proposed NSD-DIL and the DRK are remarkably suitable for real-time embedded applications of deblurring and ISR tasks.

### v2

@@ 5 Conclusion
In this work, our goal is to design a deblur model to restore sharp images from blurry inputs for practical applications. Therefore, we proposed $D^{3}$, a computationally efficient deblur model and the first DL-based image data-independent model that involves a restoration process rather than learning a mapping from distorted to sharp image distribution. Our proposed deblur framework can be directly extended to the ISR task as well. The superior performance of the proposed method can be attributed to two main reasons: the DIL objective and the RKG dataset. This work demonstrates tremendous hope for improving the deblur and ISR capability without the need for image datasets (i.e. both supervised and unsupervised ISR datasets). Our $D^{3}$ method paves the path to have a deeper look into the learning and understanding of the inverse degradation kernel since the explicit representation i.e. DRK, in a matrix form is made possible. Additionally, we have demonstrated the deblurring task of replacing the deep model with a simple convolution operation with the computed DRK, which doesn’t require high-performance computing resources. Despite being a very lightweight deblurring model, the proposed method’s performance is superior to the SotA works. The experimental results and computational performance comparisons with the SotA indicate that the proposed $D^{3}$ and the DRK are remarkably suitable for practical embedded applications of deblurring and ISR tasks.